OpenStudy (lovelycharm):

solve each system using substitution y = 2 x + 5 y = 6 x + 1

OpenStudy (anonymous):

substitute y=2x+5 in the 2nd. eqt. what do u get?

OpenStudy (lovelycharm):

how do i do it @ghass1978

OpenStudy (anonymous):

y=6x+1 now put instead of y 2x+5 =>2x+5=6x+1 can you solve this one for x?

OpenStudy (lovelycharm):

how i solve for x

OpenStudy (anonymous):

yo need to isolate x: 2x-6x=1-5 =>-4x=-4=>x=1

OpenStudy (lovelycharm):

what i suppose to do next

OpenStudy (anonymous):

find y from eitehr of the equations by substituting 1 instead of x

OpenStudy (lovelycharm):

so how i do it

OpenStudy (anonymous):

y=6x+1 put 1 instead of x

OpenStudy (lovelycharm):

so i add or subtract

OpenStudy (anonymous):

y=6(1)+1=???

OpenStudy (lovelycharm):

so it 7

OpenStudy (anonymous):

yeah

OpenStudy (lovelycharm):

thank u